Romero issues $3,400 of 10%, 10 year bonds dated January 1, 2019, that pay interest semiannually on June 30 and December 31. The bonds are issued at a price of $4,192,932.

1. Prepare the January 1 journal entry to record the bonds issuance.

2. For each semiannual period, compute (a) the cash payment, (b) the straight line discount amortization, and (c) the bond interest expense.

3. Determine the total bond interest expense to be recognized over the bonds' life.

4. Prepare the first two years of a straight line amortization table.

5. Prepare the journal entries to record the first two interest payments.
Answer #1

1.

Date Account Titles Debit Credit
Jan-01 Cash $   41,92,932
       Bonds Payable $   34,00,000
       Premium on Bonds Payable $     7,92,932

2.
(a) Cash Payment = $3,400,000 x 10% x 1/2 = $170000
(b) Straight line Premium Amortization = $792932/20 = $39646.60
(c) Bond Interest Expense = $170000-39646.60 = $130353.40

3. Total Bond interest expense = $3400000 x 10% x 10 - $792932 = $2607068

4.

Period Cash Payment Premium Amortization Interest Expense Carrying Value
0 $ 41,92,932.00
1 $ 1,70,000.00 $   39,646.60 $   1,30,353.40 $ 41,53,285.40
2 $ 1,70,000.00 $   39,646.60 $   1,30,353.40 $ 41,13,638.80
3 $ 1,70,000.00 $   39,646.60 $   1,30,353.40 $ 40,73,992.20
4 $ 1,70,000.00 $   39,646.60 $   1,30,353.40 $ 40,34,345.60

5.

Date Account Titles Debit Credit
Jun-30 Interest Expense $ 1,30,353.40
Premium on Bonds Payable $     39,646.60
      Cash $   1,70,000.00
Dec-31 Interest Expense $ 1,30,353.40
Premium on Bonds Payable $     39,646.60
      Cash $   1,70,000.00